find the area of the region bounded by the lines y=x y=3x and x+y=4 using double integral


1
Expert's answer
2021-01-26T19:29:30-0500

The intersection between the lines "y=x" and "y=4-x" is,


"x=4-x"


"2x=4"


"x=2"


So, the point of intersection is "(2,2)"


The intersection between the lines "y=3x" and "y=4-x" is,


"3x=4-x"


"4x=4"


"x=1"


So, the point of intersection is "(1,3)"


The sketch of the bounded region is as shown in the figure below:





The area of the region bounded by the lines is evaluated as,


Area"(A)=\iint_{R_1}+\iint_{R_2}"


"=\int_{x=0}^{1}\int_{y=x}^{3x}dydx+\int_{x=1}^{2}\int_{y=x}^{4-x}dydx"


"=\int_{x=0}^{1}[y]_{y=x}^{3x}dx+\int_{x=1}^{2}[y]_{y=x}^{4-x}dx"


"=\int_{x=0}^{1}(3x-x)dx+\int_{x=1}^{2}(4-x-x)dx"


"=2\int_{x=0}^{1}xdx+\int_{x=1}^{2}(4-2x)dx"


"=2[\frac{x^2}{2}]_{x=0}^{1}+[4x-2(\frac{x^2}{2})]_{x=1}^{2}"


"=1+4(2-1)-(2^2-1^2)"


"=1+4-3"


"=1+1"


"=2"


Therefore, the area of the bounded region is Area"(A)=2" square units.